OpenStudy (anonymous):

Hey can you please help ill give a medal i just dont get this question. Thanks Solve for x. Show your work. Quadrilateral A, B, C, D, showing angle measures as follows: Angle A is the quantity twenty-five times x degrees, Angle B is eighty degrees, Angle C is seventy-five degrees, and Angle D is fifty-five degrees.

OpenStudy (anonymous):

@winry can you help me

OpenStudy (jdoe0001):

got picture?

OpenStudy (anonymous):

Yes sir

OpenStudy (anonymous):

Ill give you a medal now thanks for helping me. I trust you'll help me.

OpenStudy (anonymous):

\[360=80+75+55+25x\]\[25x=150\]\[x=\frac{150}{25}\]\[x=6\]

OpenStudy (jdoe0001):

a quadrilateral's internal angles sum is 360 degrees so one could say that \(\bf \measuredangle A+\measuredangle B+\measuredangle C+\measuredangle D = {\color{red}{ 360^o}} \\ \quad \\ 25x+80+55+75=360\qquad \textit{solve for "x"}\)

OpenStudy (anonymous):

How did you get different answers

OpenStudy (anonymous):

Not to be rude.

OpenStudy (anonymous):

Thanks guy I appreciate this very much.